# Mixing High Harmonic Generation and X-ray Second Harmonic Generation

### v1

@@ 3 Results and discussion
Realizing HHG-XSHG experimentally requires intense x-ray pulses, which can be generated by advanced XFEL lasers. In these experiments, macroscopic contributions can be crucial for achieving detectable signals, with phase matching being a key factor for efficient harmonic generation. Minimizing the phase mismatch term is essential for coherent signal accumulation, and techniques developed for laser-driven HHG can be adapted to optimize phase matching in HHG-XSHG [36].
@@
In a HHG-XSHG experiment, several competing processes can impact the signal. One significant process is ionization from valence shells, where intense optical laser pulses ionize valence electrons, shifting the two-photon resonance and introducing additional radiation pathways. This ionization can be mitigated by using lower-intensity optical pulses while still maintaining interaction with the x-ray-excited state. Moreover, Auger-Meitner decay can deplete the excited-state population, competing with HHG-XSHG. This decay occurs within a few femtoseconds (e.g., 2 to 3 fs for neon) [37], while the rescattering process can take less than one optical cycle (approximately 2.67 fs for 800 nm) or even shorter for short trajectories (less than 1.3 fs) [29]. Consequently, Auger-Meitner decay filters out contributions from long-trajectory rescattering. Additionally, fluorescence from electrons returning to lower energy states introduces background noise near 846 eV (corresponding to the 3p to 1s transition), but this is spectrally distinct from the SHG signal.
@@
It is crucial to highlight that the proposed laser-assisted process fundamentally differs from previously demonstrated four-wave mixing of x-rays, even in the presence of optical lasers [38, 39, 40, 41, 42, 43, 44]. In our HHG-XSHG process, the additional optical laser plays a pivotal role by inducing strong-field excitation or facilitating tunneling ionization of x-ray-excited electrons. This leads to a highly nonlinear response to the optical field, which is distinctly different from the mechanisms driving traditional four-wave mixing. Additionally, HHG-XSHG distinguishes itself from previous XUV-assisted HHG approaches [45, 46, 47, 48]. In XUV-assisted HHG, ionization typically occurs from the valence shell, with XUV pulses primarily serving to enhance conversion efficiency and extend the cutoff energy within the XUV spectral range through single-photon excitation or ionization. In contrast, our method leverages two-photon excitation and ionization of core electrons, enabling background-free detection, enhancing efficiency, and extending the emitted radiation into the x-ray spectral domain. This advancement provides a robust platform for probing ultrafast dynamics of core electrons in atoms and molecules.
@@
Furthermore, broadband HHG-XSHG can generate ultrashort pulses with durations on the order of a few attoseconds, thanks to its broad spectral bandwidth, which is determined by the intensity and wavelength of the driving laser. By employing a mid-infrared driving laser, the bandwidth of the HHG-XSHG can be extended to over 1 keV, corresponding to a Fourier transform-limited pulse duration of approximately 1.8 attoseconds [49]. These ultrashort pulses are essential for probing valence and core electron dynamics on their intrinsic timescales, thereby opening new avenues for time-resolved studies of quantum processes and ultrafast phenomena in a variety of systems.
@@
4 Conclusion and outlook
In conclusion, we have theoretically demonstrated the feasibility of HHG-XSHG through detailed simulations on gas-phase atoms. Our results reveal a significant enhancement of the HHG-XSHG signal via two-photon excitation resonance, underscoring the crucial importance of tuning the x-ray photon energy to match core-level transitions. This approach enables efficient HHG-XSHG generation in gases. The HHG-XSHG mechanism combines the laser-driven HHG with x-ray SHG, facilitating the application of attosecond techniques developed in HHG to the x-ray regime, alongside XFEL-driven nonlinear x-ray techniques. This opens new possibilities for generating element-specific attosecond x-ray pulses and probing ultrafast core-electron dynamics and nonlinear interactions with high temporal and spectral resolution. HHG-XSHG offers a powerful tool for studying electronic structures, excited-state dynamics, and symmetry properties, positioning it as a transformative technique in attosecond science and nonlinear x-ray science.
Acknowledgments
We gratefully acknowledge the computational resources provided by the PSI Merlin HPC cluster.

### v3

@@ III.4 Further discussion
Realizing HHG-XSHG experimentally requires intense x-ray pulses, which can be generated by advanced XFEL lasers with their short pulse durations, intense pulse energies and photon energy tunability Prat et al. (2020); Habib et al. (2023); Mcneil and Thompson (2010); Prat et al. (2023); Sobolev et al. (2020); Mcneil and Thompson (2010); Shwartz et al. (2014); Lam et al. (2018). In these experiments, macroscopic contributions can be crucial for achieving detectable signals, with phase matching being a key factor for efficient harmonic generation. Minimizing the phase mismatch term is essential for coherent signal accumulation, and techniques developed for laser-driven HHG can be adapted to optimize phase matching in HHG-XSHG Popmintchev et al. (2009).
@@
In a HHG-XSHG experiment, several competing processes can impact the signal. One significant process is ionization from valence shells, where intense optical laser pulses ionize valence electrons, shifting the two-photon resonance and introducing additional radiation pathways. This ionization can be mitigated by using lower-intensity optical pulses while still maintaining interaction with the x-ray-excited state. Moreover, Auger-Meitner decay can deplete the excited-state population, competing with HHG-XSHG. This decay occurs within a few femtoseconds (e.g., 2 to 3 fs for neon) Haynes et al. (2021), while the rescattering process can take less than one optical cycle (approximately 2.67 fs for 800 nm) or even shorter for short trajectories (less than 1.3 fs) Corkum (1993). Consequently, Auger-Meitner decay filters out contributions from long-trajectory rescattering. Additionally, fluorescence from electrons returning to lower energy states introduces background noise near 846 eV (corresponding to the 3p to 1s transition), but this is spectrally distinct from the SHG signal.
@@
It is crucial to highlight that the proposed laser-assisted process fundamentally differs from previously demonstrated four-wave mixing of x-rays, even in the presence of optical lasers Mukamel (2005); Rouxel et al. (2020); Morillo-Candas et al. (2024); Peters et al. (2023); Rouxel et al. (2021); Alexander et al. (2024); Glover et al. (2012). In our HHG-XSHG process, the additional optical laser plays a pivotal role by inducing strong-field excitation or facilitating tunneling ionization of x-ray-excited electrons. This leads to a highly nonlinear response to the optical field, which is distinctly different from the mechanisms driving traditional four-wave mixing. Additionally, HHG-XSHG distinguishes itself from previous XUV or x-ray-assisted HHG approaches Troß et al. (2022); Sarantseva et al. (2018); Gaarde et al. (2005); Buth et al. (2011); Tudorovskaya and Lein (2014); Buth (2015); Buth et al. (2013). In XUV-assisted HHG, ionization typically occurs from the valence shell, with XUV pulses primarily serving to enhance conversion efficiency and extend the cutoff energy within the XUV spectral range through single-photon excitation or ionization. In contrast, our method leverages two-photon excitation and ionization of core electrons, enabling background-free detection, enhancing efficiency, and extending the emitted radiation into the x-ray spectral domain. This advancement provides a robust platform for probing ultrafast dynamics of core electrons in atoms and molecules.
@@
Furthermore, broadband HHG-XSHG can generate ultrashort pulses with durations on the order of a few attoseconds, thanks to its broad spectral bandwidth, which is determined by the intensity and wavelength of the driving laser. By employing a mid-infrared driving laser, the bandwidth of the HHG-XSHG can be extended to over 1 keV, corresponding to a Fourier transform-limited pulse duration of approximately 1.8 attoseconds Popmintchev et al. (2012). These ultrashort pulses are essential for probing valence and core electron dynamics on their intrinsic timescales, thereby opening new avenues for time-resolved studies of quantum processes and ultrafast phenomena in a variety of systems.
@@
Back to the model used in the simulations, SAE approximation is widely used for solving the TDSE in strong-field and ultrafast x-ray interactions due to its computational efficiency and ability to capture essential features of HHG and related nonlinear processes Liu et al. (2020); Chen et al. (2012). However, this approximation inherently neglects multi-electron effects, including electron-electron correlations, dynamic polarization, and collective interactions, which can significantly influence the response of atomic and molecular systems. While the SAE model remains valid for describing strong-field dynamics in atoms with a well-separated core and valence structure, its applicability becomes questionable in systems where electron correlation plays a crucial role, such as in open-shell atoms, molecules, and many-electron systems with significant configuration mixing. Beyond the SAE approximation, multi-electron effects can be accounted for through methods such as time-dependent configuration interaction (TDCI) and time-dependent density functional theory (TDDFT) Bonafé et al. (2024); Guan et al. (2022); Liu et al. (2020). These approaches explicitly incorporate electron correlation, which is essential for accurately describing Auger decay, shake-up and shake-off processes, and multi-electron excitations induced by intense x-ray fields. Future studies will explore these advanced techniques to improve the accuracy of our model, particularly in systems where strong correlation effects may alter the HHG-XSHG response. Additionally, including electron correlation will provide deeper insights into the role of collective excitations and correlation-driven resonances in nonlinear x-ray interactions, further advancing our understanding of attosecond x-ray spectroscopy in complex systems Jin et al. (2025).
IV Conclusion and outlook
In conclusion, we have theoretically demonstrated the feasibility of HHG-XSHG through detailed simulations on gas-phase atoms. Our results reveal a significant enhancement of the HHG-XSHG signal via two-photon excitation resonance, underscoring the crucial importance of tuning the x-ray photon energy to match core-level transitions. This approach enables efficient HHG-XSHG generation in gases. The HHG-XSHG mechanism combines the strong-field-driven HHG with x-ray SHG, facilitating the application of attosecond techniques developed in HHG to the x-ray regime, alongside XFEL-driven nonlinear x-ray techniques. This opens new possibilities for generating element-specific attosecond x-ray pulses and probing ultrafast core-electron dynamics and nonlinear interactions with high temporal and spectral resolution. HHG-XSHG offers a powerful tool for studying electronic structures, excited-state dynamics, and symmetry properties, positioning it as a transformative technique in attosecond science and nonlinear x-ray science.